OpenStudy (anonymous):

Please Help!!! Find an equation of the line that has a slope of 6 and a y intercept of -9. Write your answer in the form y=mx+b . How do I solve this for y? y=?

OpenStudy (anonymous):

see slop intercept form od a line y = mx + c

OpenStudy (anonymous):

m is slope nd c is intercept........put those values

OpenStudy (anonymous):

@A.Avinash_Goutham So -9=6x + b ? How would I solve this now to get y?

OpenStudy (anonymous):

noo............c is y intercept

OpenStudy (anonymous):

So the answer is c? I'm lost @A.Avinash_Goutham

OpenStudy (anonymous):

noo y = 6x - 9

OpenStudy (anonymous):

take ur time to solve things........dont give up soo easily

OpenStudy (anonymous):

So the answer is y=6x-9 ? So I'm not solving just looking for an equation? Is that what I was suppose to do? @A.Avinash_Goutham

OpenStudy (sriram):

ur question is to look for the equation

OpenStudy (anonymous):

That explains why I got it wrong lol I was trying to solve it... ;( Losing my mind I don't get slope !!!!! @Sriram @A.Avinash_Goutham

OpenStudy (sriram):

look u have to write the eq which satisfies the given conditions y= mx +c now m tells how y varies when x varies m is called the slope (in ur case it is given 6) and c is the value of y when x is 0 meaning the y intercept (in ur case tis -9) hence just write the eq again y=mx +c y= 6x + (-9) y=6x-9